OpenStudy (anonymous):

Simplify the number using the imaginary unit i. Sqrt -28

OpenStudy (anonymous):

You can write : \[\sqrt{-28} = \sqrt{-1 \times 28} = \sqrt{-1} \times \sqrt{28}\] And : \(\sqrt{-1} = i\) Plug in the above expression and tell me what did you get ??

OpenStudy (mayankdevnani):

\[\sqrt{-28}=7\times2\times2\] \[\sqrt{-28}=-2\sqrt{7}\] OR \[\sqrt{-28}=2\sqrt{-7}\]